6 boys and 6 girls sit in a row at random. Find the probability that
1) The six girls sit together
2) The boys and girls sit alternatively.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
P(group of girls)
= 7!6! / (6+6)!
= 7!6! / 12!
= 6*5*4*3*2*1 / 12*11*10*9*8
= 1/132
P(alternately)
= 2 * 6!6! / 12!
= 1/462
Thanks and Regards
Shaik Aasif
askIITians faculty